News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Thursday 12 August 1999 7.00am EST

The following is an overview of the events that occurred over the last 24 hours.

Successfully opened the HRMA Aft Contamination Cover at approximately 11:00pm EDT on 8/11/99.

Completed Earth Scan #24, however, the data was not retrieved due to the use of an incorrect telemetry format. However, indications are that our current attitude reference is sufficiently accurate that without this update the upcoming activities will not be impacted.

Activated HRC to assess its ability to achieve operational high voltage with the current optical bench pressure. Results of this test appear to indicate that pressure is still at a value that will not permit high voltage operation of the HRC. No more attempts to operate HRC high voltage are now planned until after "first light" and it is translated into the telescope focal plane.

Began ACA dark current calibration.

Began Flight Contamination Monitor (FCM) calibration using ACIS.

Activities for next 24 hours.
  • Continue ACIS FCM calibration.
  • Open Chandra Sun Shade Door.
  • Begin PCAD calibration using ACA.






Return to Status Reports